Which of the following statements is/are true? Explain why the false one is false.
A. If u = (u1; u2) with ui 6= 0 for some 1  i  2; then the unit vector in the direction of u is
p 1
u21
+u22
(u1; u2) : (2)
B. u = (1; 2; 3) is the unit vector in the direction of u: (
1
Expert's answer
2020-04-15T11:11:02-0400

(A)True because unit vector in the direction of u is "p_1=\frac{u(u_1,u_2)}{\sqrt{u_1^2+u_2^2}}"

(B)False because u(1,2,3) is not the unit vector in the direction of (1,2,3), the vector v"(\frac{1}{\sqrt{14}},\frac{2}{\sqrt{14}},\frac{3}{\sqrt{14}})" is the unit vector in the direction of (1,2,3).